OpenStudy (anonymous):

A cell phone company orders 600 new phones from a manufacturer. If the probability of a phone being defective is 3.5%, predict how many of the phones are likely to be defective. Round to the nearest whole number... A.) 210 phones.. B.) 18 phones C.)21 phones D.)26 phones

OpenStudy (kropot72):

Find 3.5% of 600:\[600\times \frac{3.5}{100}=you\ can\ calculate\]

OpenStudy (anonymous):

Probability of a defective phone is 3.5%, as a fraction that is 7/200. So it follows: 600 * 7/200 = 4200 / 200 = 21--> Therefore, there is about 21 phones that will be defective.

OpenStudy (anonymous):

thanks